Find two positive numbers whose product is 25 and whose sum is a minimum

Find: Smaller number, Larger number
Answer #1

xy=25
x=25/y
x+y=Minimum
25/y+y=minimum
take the derivative of the minimum function to get:
-(25/y2)+1

set equal to 0

-(25/y2)+1=0

solve for y

then plug that back in to the original equation to solve for x

answered by: dda